# Latent Policy Steering with Embodiment-Agnostic Pretrained World Models

### v2

@@ Abstract
The performance of learned robot visuomotor policies is heavily dependent on the size and quality of the training dataset. While an increasing amount of robot or human data is publicly available, embodiment gaps and differing action spaces can make leveraging these data challenging. In this work, we aim to learn visuomotor policies in a low-data regime by pretraining a model across embodiments (e.g., humans and multiple robots). Our approach pretrains a World Model (WM) with embodiment-agnostic action representations based on optical flow on easily-collected or readily available multi-embodiment data. This WM is finetuned on a small dataset collected on the target robot-embodiment ( data collection << 1 hour ). To leverage this WM, we develop Latent Policy Steering (LPS), a technique that rolls out action candidates and identifies the best action during inference based on a learned value function robust to distribution shifts. In the real-world experiments, our method achieves over 50% relative improvement with 30-50 target-embodiment demonstrations, and over 25% relative improvement with 50-100 demonstrations, compared to a pure imitation learning baseline.
@@ I INTRODUCTION
Imitation learning through Behavior Cloning (BC) is a widely adopted paradigm to acquire visuomotor policies for robots [1][2][3]. To achieve high task success, sufficient expert demonstrations must be collected, which is a time-consuming process. Furthermore, the data collected are often specific to a robot, a task, or an environment, and the data collection process may need to be repeated for different embodiments or different environments. Thanks to the progress in collecting large datasets [4] across different robots and environments, progress has been made to build generalist robot policies [5][6] via cross-embodiment training. However, these models sometimes do not generalize to new tasks with satisfactory results. To finetune these models for better performance, a considerable amount of data is required, given the large model size used to learn from large datasets [1][3][7]. For example, $\pi 0$ [7] requires 5-10 extra hours of fine-tuning data to achieve high success rates on new tasks.
@@
According to historical Machine Learning research, a model trained on large and diverse datasets across multiple tasks will produce general representations that are transferable to new tasks with only a small amount of data, known as pretraining [6][8]. However, pretraining a robot model is challenging: each robot in the dataset has embodiment-specific information (proprioception, actions). Thus, a pretraining dataset with different embodiments makes the pretrained representation heavily dependent on the embodiments included in the dataset and may generalize poorly to new embodiments.
To tackle this unique challenge, we choose to pretrain a robot World Model (WM) with an embodiment-agnostic action representation: optical flow (Fig. 1). Since optical flow captures visual motion in the pixel space, it is less dependent on the specific embodiment and allows different embodiments to share the same action space. Instead of a policy, we choose to pretrain a WM with optical flow because a WM can leverage suboptimal data. During inference, such a WM will improve a policy via policy steering [9, 10, 11] to improve the performance of the policy.
Figure 1: World Model (WM) pretraining with optical flow as an embodiment-agnostic action representation. Thanks to optical flow, we can integrate data from multiple types of embodiments (robots, humans), with a shared action space. A WM is chosen for pretraining since it can leverage suboptimal data. We encode every optical flow to a vector input of the WM, and train the encoder end-to-end with the WM.
To leverage the pretrained WM, we propose Latent Policy Steering (LPS), a technique to improve a policy with the WM during inference, motivated by prior work on using a WM for planning and control [12][13]. In that work, the authors simulate the action plan sampled from the policy with the WM by predicting future states and comparing them to a goal image in the latent space. Unfortunately, that technique becomes less effective when the task is long-horizon and the desired outcome (goal image) is beyond the reach of the policy and the WM. Our proposed technique is effective regardless of the task horizon by making one key observation: since every demonstration collected for Behavior Cloning is an expert demonstration that accomplishes the task, every state in the demonstration becomes a goal state. Thus, one can distill the goal-state-to-state comparisons conducted during inference into a state-based value function. Such a value function steers a policy close to the dataset (pessimism [14][15][16][17]), and can approximate the results beyond the horizon. Since behavior cloning methods are vulnerable to distribution shifts, we learn a robust value function on both in-distribution and out-of-distribution data. During inference, the robust value function steers a policy toward success by visiting states with higher rewards and closer to the dataset distribution.
@@
  • We propose Latent Policy Steering (LPS) to improve the policy during inference with a WM and a robust value function. To avoid distribution shift, we learn the robust value function on in-distribution and out-of-distribution data. During inference, the LPS steers the policy towards states that have higher rewards and are closer to the dataset rather than deviating from them.
  • We demonstrate that LPS with a pretrained WM consistently improves the policy’s performance both in simulated and real-world object manipulation tasks. This shows the effectiveness of improving a policy in small data scenarios by pretraining WMs on abundant task-relevant data from other embodiments.
@@ II-A Robot Learning from Diverse Data
Collecting large amounts of robot data for a specific task can be challenging and time-consuming. A common strategy to overcome this limitation is to leverage data from other sources, including online robot datasets with multiple types of robots and even human videos. Prior works [18][19][20][21] train a visual encoder using both human and robot video data to avoid learning the policy network from scratch. However, the pretrained encoder only provides visual representations instead of directly learning to make decisions. Thus, to learn full policies across multiple robot embodiments, recent works have also looked at learning a single network across multiple robot embodiments by using a modular structure with a common trunk and various heads for different action spaces [5][22]. Another common approach is the vision-language-action model (VLA) approach, where a pretrained vision-language model is finetuned with robot data to act as a robot policy [23][7][24]. However, due to their large size, both these approaches can be expensive to finetune to improve performance on a specific task, and sometimes suffer from slow inference. We compare our approach against a finetuned HPT [22] as a representative example of these types of models.
@@ II-B Policy Steering & Planning with World Models
Recurrent-state-space-based world models (RSSM) have become an increasingly common approach to model environment dynamics and transition functions in robotics. Popularized by Hafner et. al [27][28], these models consist of an encoder, which maps visual observations to a latent state representation, a latent transition function which predicts future latent states, and a decoder which is used primarily during training to propagate gradients. Once trained, these world models can be utilized in a variety of ways. One approach is to learn actor and critic models from the same training data to generate action sequences, roll them out over time, and evaluate them with the critic [28]. Another approach is to encode a goal image in the latent space and then use the similarity between planned states and the goal image as a value function to select the best sequence of actions. These action sequences are commonly optimized using gradient-free optimizers like the Cross-Entropy Method (CEM), but gradient-based optimization can also be used [12][13][28].
@@ II-C Offline & Inverse Reinforcement Learning
We propose a value function that favors states in the distribution of the training data, which is related to the pessimism used in offline RL [14][15][16][17] and the concept of Inverse RL [29]. In offline RL, a model ( e.g., value functions, dynamics models) only has access to the state-action pairs in an offline dataset. Therefore, predictions for state-action pairs outside of the training data distribution could be unreliable. For example, a value function could be overly optimistic when queried with out-of-distribution state-action pairs, leading to extrapolation errors [14].
@@ II-D Optical Flow as a Representation in Robotics
Figure 2: Overview of the method. The figure illustrates all the phases of our method, including WM pretraining, finetuning, and Latent Policy Steering (LPS). After pretraining the WM with optic flow actions, the second phase involves collecting a small robot dataset to learn a policy from scratch, finetuning the WM, and learn a robust value function based on the WM. During inference, the WM and value function are used to select the best candidate.
@@ III Preliminaries
We consider robot control in a Partially Observable Markov Decision Process (POMDP) setting, with a latent state space learned by a neural network. A POMDP can be described by a tuple $(\Omega,\mathcal{X,S,A,R,P})$, which consists of observations $x\in\mathcal{X}$, conditional observation probabilities $\Omega(x^{\prime}|s,a)$, states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$, reward function $r=\mathcal{R}(s,a)$, and transition function $P(s^{\prime}|s,a)$. At the timestep $t\in[1,2,...,T]$, the observation, state, action and reward are denoted as $o_{t},s_{t},a_{t},r_{t}$. Given discount factor $\gamma$, the expected future discounted reward is defined as: $E[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. A dataset consists of sequences of ($x_{t},a_{t},r_{t}$) where $r_{t}$ is a binary reward indicating task successes. Given a horizon length $h$, policy steering seeks the best action plan $\textbf{a}_{t}^{*}=[a^{*}_{t},a^{*}_{t+1},...,a^{*}_{t+h}]$ such that: $\textbf{a}_{t}^{*}=$ argmax a∈A $\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\mathcal{R}(s,a)$.
We assume a robot has access to a small dataset of demonstrations for a task on its embodiment (referred to as the target embodiment) and a larger dataset of demonstrations involving similar objects or for a similar task on non-target embodiments, such as other robots or humans.
Figure 3: Optical flow as an embodiment-agnostic action representation. We observe that motions captured by optical flow across embodiments are similar in visual space. By using optical flow as an action representation, we remove WM’s dependency on specific embodiments, allowing the pretrained model to efficiently leverage data from multiple embodiments.
@@ IV Methods
In this section, we first explain our motivation for choosing optical flow as an embodiment-agnostic action representation for pretraining a WM. Then, we describe the proposed technique called Latent Policy Steering (LPS) to leverage a WM to improve the performance of a policy during inference.
@@ IV-A Flow-as-Action: Embodiment-Agnostic World Modeling
A robot WM that can be easily adapted to a new embodiment should be less dependent on embodiment-specific information during pretraining, such as proprioception and actions. Specifically, to model task dynamics, WMs should use an embodiment-agnostic action representation rather than a raw robot action as input (embodiment-specific). Unlike prior work, such as HPT [22], which learns generalizable representations on the fly with end-to-end training, we propose to replace the raw action from different embodiments with an existing embodiment-agnostic action representation that is easy to compute with the off-the-shelf tools: optical flow. One of our key observations is that different robots’ visual motions have similar patterns when they execute similar skills ( e.g., pick up an object), as shown in Fig. 3. Therefore, we leverage optical flow as an embodiment-agnostic action representation for WM’s pretraining.
We precompute optical flow for each episode in the pre-training dataset. A Convolution Neural Network (CNN) with a Multi-Layer Perceptron (MLP) is used to encode the optical flow and project the encoding to a vector with a small dimension (e.g., a 7-dimensional vector). During pretraining, this encoder is optimized with the WM fully end-to-end. When finetuning the WM for a target embodiment, the optical flow encoder is discarded, and the WM takes raw robot actions without any extra encoding. While this does create a mismatch between latent flow representations during pretraining and robot actions during finetuning, we find the finetuning process is easily able to compensate for this.
IV-B Latent Policy Steering with a world model
We propose Latent Policy Steering (LPS) to improve the policy’s performance during inference. Similar to prior work [9][11], policy steering is conducted by sampling $B$ number of candidate plans from the policy (phase 3 in Fig. 2), selecting the best action plan according to the value function, and executing the plan in open-loop. Differently, our value function is more robust to distribution shifts and also guides the policy to stay close to the dataset. The following paragraphs describe our reasoning for each design.
Robustness
A vanilla value function learns to estimate future rewards and selects the best action accordingly [36]. However, in a small dataset, successful demonstrations only cover a limited set of states, and a policy may not visit the same set of states during inference. The resulting distribution shifts will make a vanilla value function not reliable for policy steering. To make our value function robust, we train the value function on 1) states in the dataset ($s_{t}$ expert states), and 2) states likely to be visited by the policy ($s^{\prime}_{t}$ non-expert states) that could be out of the training data distribution. The expert states are the results of unroll the WM with the actions in the dataset (line 7, Alg. 1). The non-expert states are computed by 1) querying the policy with the current observation to predict actions (line 5, Alg. 1), and 2) unrolling the WM with predicted actions (line 7, Alg. 1).
Pessimism
Since we are in an offline environment, we would like to stay close to the dataset instead of deviating from it, similar to the pessimism in offline RL [14][15][16][17]. We assume that the expert and non-expert states generated from the same observation have a 1-to-1 correspondence (line 6-7, Alg. 1). Thus, the non-expert state also shares the same environmental binary rewards as its corresponding expert state. We create a penalization for non-expert states that are deviating from the expert states based on the cosine distance between them (line 8, Alg. 1). Lastly, we create a lambda return target for the value function, similar to the latest Dreamer [28], shown from lines 9-10, Alg. 1.
Algorithm 1 Value function training for LPS.
1:Requires: Robot dataset $D=[(x_{1:T},a_{1:T},r_{1:T})_{n}|n=[1,...,N]]$ with observations, actions, and binary
2:rewards, the frozen world model $\mathcal{W}$ and policy $\pi$.
3:Initialize: value function $\mathcal{V}$, horizon $h$, discount $\gamma$, and $\lambda$ for return computation.
4:Samples $(x_{t:t+h},a_{t:t+h},r_{t:t+h})$ from $D$.
5:Sample a predicted action plan: $a^{\prime}_{t:t+h}\sim\pi(x_{t})$.

### v3

@@ Abstract
The performance of learned robot visuomotor policies is heavily dependent on the size and quality of the training dataset. Although large-scale robot and human datasets are increasingly available, embodiment gaps and mismatched action spaces make them difficult to leverage. Our main insight is that skills performed across different embodiments produce visual similarities in motions that can be captured using off-the-shelf action representations such as optical flow. Moreover, World Models (WMs) can leverage sub-optimal data since they focus on modeling dynamics.
In this work, we aim to improve visuomotor policies in low-data regimes by first pretraining a WM using optical flow as an embodiment-agnostic action representation to leverage accessible or easily collected data from multiple embodiments (robots, humans). Given a small set of demonstrations on a target embodiment, we finetune the WM on this data to better align the WM predictions, train a base policy, and learn a robust value function. Using our finetuned WM and value function, our approach evaluates action candidates from the base policy and selects the best one to improve performance. Our approach, which we term Latent Policy Steering (LPS), improves behavior-cloned policies by 10.6% on average across four Robomimic tasks, even though most of the pretraining data comes from the real world. In the real-world experiments, LPS achieves larger gains: 70% relative improvement with 30-50 target-embodiment demonstrations, and 44% relative improvement with 60-100 demonstrations, compared to a behavior-cloned baseline.
@@ I INTRODUCTION
Imitation learning through Behavior Cloning (BC) is a widely adopted paradigm to acquire visuomotor policies for robots [5][8][31]. To achieve high task success, sufficient expert demonstrations must be collected, which is a time-consuming process. Furthermore, the data collected is often specific to a robot, a task, or an environment, and the data collection process may need to be repeated for different embodiments or different environments. By collecting large datasets across different robots and environments [34], progress has been made toward building generalist robot policies via cross-embodiment training [32][14]. However, these models sometimes do not generalize to new tasks with satisfactory results. To finetune these models for better performance, a considerable amount of data is required, given the large model size used to learn from large datasets [5][31][3]. For example, $\pi 0$ [3] requires 5-10 extra hours of fine-tuning data to achieve high success rates on new tasks.
@@
In typical LLM/VLM practice, a model trained on large and diverse datasets across multiple tasks will produce general representations that are transferable to new tasks with only a small amount of data, known as pretraining [14][15]. However, pretraining a robot model is challenging: each robot in the dataset has embodiment-specific information (proprioception, actions). Thus, a pretraining dataset with different embodiments makes the pretrained representation heavily dependent on the embodiments included in the dataset and may generalize poorly to new embodiments.
To tackle this challenge, prior works attempt to extract shared action across embodiments through large-scale unsupervised learning [30][6]. However, we propose a more straightforward solution given our key observation: skills performed across different embodiments produce visually similar motion, which can be captured by off-the-shelf tools such as optical flow. Optical flow allows different embodiments to share the same action space, as shown in Fig. 1. We learn an image-based WM (Dreamer [12][13]) with encoded optical flow as action (middle of the Fig. LABEL:fig:motivation1). To encode optical flow, we train a convolution-based encoder end-to-end with the WM to produce compact encodings that suppress motion-irrelevant information (e.g., noise, morphology differences across embodiments). During pretraining, we choose to learn a WM instead of a policy, since WM learns dynamics and is more sample-efficient than a policy [12][13].
For a given target embodiment and task, we collect a small expert dataset to learn a diffusion policy [8] from scratch. The WM is finetuned on the target dataset with robot actions instead of optical flow, and a value function is extracted. During the inference, the WM and value function will evaluate multiple action plans sampled from the policy, and select the best action plan for execution (bottom right of the Fig. LABEL:fig:motivation1). This process is often referred to as Policy Steering [25][36][37].
Prior work on Policy Steering [25][36][37] has largely concerned steering the policy toward the user’s desired outcome. In our implementation, we address a prevalent problem for sequential decision-making: distribution shift. We propose Latent Policy Steering (LPS): LPS learns a robust value function to avoid distribution shift by simulating the shift in the latent state space during finetuning. Specifically, we use the world model in conjunction with the base policy to simulate states the policy is likely to visit during inference and train our value function across these simulated states to penalize distribution shifts from the expert data. As a result, the LPS’s value function is robust to inference-time distribution shift as it has foreseen it during finetuning.
Figure 1: Optical flow as action. We observe that motions captured by optical flow across embodiments are similar in visual space. For example, picking up a cup will generate similar flow vectors in both the agent and the cup, regardless of agent embodiment. By using optical flow as an action representation, we reduce dependency on specific embodiments, allowing the pretrained model to more easily adapt to a new embodiment during finetuning.
@@
  • We propose Latent Policy Steering (LPS) to align the embodiment-agnostic WM to a target-embodiment policy. To mitigate inference-time distribution shift, LPS learns a robust value function in the WM’s latent space by training on both expert states and states policy is likely to visit during inference, and steers the policy back towards the expert state distribution via latent state similarities between them.
  • We demonstrate that LPS with a pretrained WM significantly improves the policy’s performance on long-horizon manipulation in simulations, and manipulation tasks in the real-world involving tool-use and deformable objects. Our results show the effectiveness of pretraining embodiment-agnostic WMs on abundant data from other embodiments to enhance a policy in low-data scenarios.
@@ II-A Robot Learning from Diverse Data
Collecting large amounts of robot data for a specific task can be challenging and time-consuming. A common strategy to overcome this limitation is to leverage data from other sources, including online robot datasets with multiple types of robots and even human videos. Prior works [16][24][21][28] train a visual encoder using both human and robot video data to avoid learning the policy network from scratch. However, the pretrained encoder only provides visual representations instead of directly learning to make decisions. Thus, to learn full policies across multiple robot embodiments, recent works have also looked at learning a single network across multiple robot embodiments by using a modular structure with a common trunk and various heads for different action spaces [32][35]. Another common approach is the vision-language-action model (VLA) approach, where a pretrained vision-language model is finetuned with robot data to act as a robot policy [4][3][18]. However, due to their large size, both these approaches can be expensive to fine-tune to improve performance on a specific task, and sometimes suffer from slow inference. We compare our approach against a finetuned HPT [35] as a representative example of these types of models.
@@ II-B Policy Steering & Planning with World Models
Recurrent-state-space-based world models (RSSM) have become an increasingly common approach to model environment dynamics and transition functions in robotics. Popularized by Hafner et al. [12][13], these models consist of an encoder, which maps visual observations to a latent state representation, a latent transition function which predicts future latent states, and a decoder which is used primarily during training to propagate gradients. Once trained, these world models can be utilized in a variety of ways. One approach is to learn actor and critic models from the same training data to generate action sequences, roll them out over time, and evaluate them with the critic [13]. Another approach is to encode a goal image in the latent space and then use the similarity between planned states and the goal image as a value function to select the best sequence of actions. These action sequences are commonly optimized using gradient-free optimizers like the Cross-Entropy Method (CEM), but gradient-based optimization can also be used [41][23][13].
@@ II-C Offline & Inverse Reinforcement Learning
We propose a value function that favors states in the distribution of the training data, which is related to the pessimism used in offline RL [10][38][17][19], and the concept of Inverse RL [26]. In offline RL, a model ( e.g., value functions, dynamics models) only has access to the state-action pairs in an offline dataset. Therefore, predictions for state-action pairs outside of the training data distribution could be unreliable. For example, a value function could be overly optimistic when queried with out-of-distribution state-action pairs, leading to extrapolation errors [10].
@@ III Preliminaries
We consider robot control in a Partially Observable Markov Decision Process (POMDP) setting, with a latent state space learned by a neural network. A POMDP can be described by a tuple $(\Omega,\mathcal{X,S,A,R,P})$, which consists of observations $x\in\mathcal{X}$, conditional observation probabilities $\Omega(x^{\prime}|s,a)$, states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$, reward function $r=\mathcal{R}(s,a)$, and transition function $P(s^{\prime}|s,a)$. At the timestep $t\in[1,2,...,T]$, the observation, state, action, and reward are denoted as $x_{t},s_{t},a_{t},r_{t}$. Given discount factor $\gamma$, the expected future discounted reward from state $s_{t}$ is defined as: $E[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$, and can be estimated by a state-value function $\mathcal{V}(s_{t})$. A dataset consists of sequences of ($x_{t},a_{t},r_{t}$) where $r_{t}$ is a binary reward indicating task successes. In this work, we consider two such datasets: A small dataset $\mathcal{E}$ of expert demonstrations on the target robot embodiment, and a larger cross-embodiment dataset $\mathcal{C}$ with action space $\mathcal{A}^{\prime}$ that may not match the action space $\mathcal{A}$ of the target embodiment and sequences that do not necessarily represent task success. Importantly, $\|\mathcal{C}\|>>\|\mathcal{E}\|$, and we would like to maximally leverage $\mathcal{C}$ in training despite its action-space mismatches and non-optimal data. Given a horizon length $h$, policy steering seeks the best action plan $\textbf{a}_{t}^{*}=[a^{*}_{t},a^{*}_{t+1},...,a^{*}_{t+h}]$ such that: $\textbf{a}_{t}^{*}=$ argmax a∈A $\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\mathcal{R}(s,a)$.
@@ IV-A Flow-as-Action: Embodiment-Agnostic World Modeling
A pretrained WM that can be easily adapted to a new embodiment should be less dependent on embodiment-specific information, such as proprioception and actions. Thus, we drop embodiment-specific information, such as proprioception, and replace embodiment-specific robot actions with an embodiment-agnostic action representation. One of our key observations is that different embodiments share similar visual motion patterns when they execute similar skills ( e.g., pick up an object), as shown in Fig. 1. Therefore, we choose optical flow, which captures visual motion and is computable via off-the-shelf tools, as the embodiment-agnostic action representation to train the WM.
Our WM adopts the objective and architecture from Dreamer v3 [13], and adds an optical flow encoder during pretraining (middle of the Fig. LABEL:fig:motivation1). The optical flow encoder is a convolution-based encoder, which predicts an $n$-dimensional vector. We set $n=\|\mathcal{A}\|$, the action space dimension of the target embodiment. Regardless of robot embodiment, the robot action space will be much smaller than the dimensionality of the optical flow image, which forces the network to extract salient features from flow and ignore noise or motion-irrelevant information (e.g., morphology differences across embodiments) when the WM is trained end-to-end on dataset $\mathcal{C}$. Setting $n$ this way also enables our simple target-embodiment finetuning procedure in the next step.
IV-B Target Embodiment Finetuning
Given our small target embodiment dataset $\mathcal{E}$, we must adapt our pretrained WM to use the target embodiment action space and steer a policy toward better actions. To this end, we first train a base policy $\pi$ with the target embodiment data (upper right of the Fig. LABEL:fig:motivation1). This policy gives us a distribution across actions for any observation $a_{t}\sim\pi(x_{t})$. To finetune the WM, we replace the optical flow encoder, which projected optical flow to an $\|\mathcal{A}\|$-dimensional vector, with normalized robot actions in this same $\|\mathcal{A}\|$-dimensional space (middle of the Fig. LABEL:fig:motivation1). The WM is trained on the target embodiment dataset using the same Dreamer v3 [13] objective. While there is naturally a distribution shift between the compressed optical flow representation and normalized robot actions, in practice, we found this approach to be both simpler and more effective than an approach that attempted to specifically align compressed optical flow with robot actions on the limited target embodiment dataset.
Finally, we also jointly train a robust value function $\mathcal{V}(s_{t})$ along with the WM, which estimates the discounted future rewards. This value function leverages the latent state representation from the WM, but its gradients do not flow back into the WM. Crucially, our value function must be able to estimate discounted future rewards across states the policy is likely to visit during inference, not just states in the expert dataset. Furthermore, the base-policy may become unstable if the robot deviates too far from states in the expert dataset, so our value function should also ensure the robot stays close to states in that expert dataset.
To train a value function with both of these qualities, we adopt the following procedure. First, given a sequence of expert data for a horizon h, we sample a predicted action plan from the policy given the same initial observation (line 4, Alg. 1), compare the resulting latent states against the latent states from the expert dataset via a similarity metric (line 6, Alg. 1), and convert this similarity into and additional reward (line 8, Alg. 1). We use cosine similarity given its prevalent usage in deep learning literature [27][7]. Thus, a robust value function is obtained by training both states likely to be visited by the policy during inference, and states from the expert dataset (line 10-11, Alg. 1), given a lambda-return objective adopted from Dreamer v3 [13].
Algorithm 1 Robust value function for LPS
1:Requires: Robot dataset $\mathcal{E}=[(x_{1:T},a_{1:T},r_{1:T})_{n}|n=[1,...,N]]$, world model $\mathcal{W}$, policy $\pi$, similarity metric simsim, lambda-return computation $r_{\lambda}$, and loss $\mathcal{L}$.
2:Initialize: the value function $\mathcal{V}$, $\mathcal{V}_{target}$, $h$, $\gamma$, and $\lambda$.
3:Samples $(x_{t:t+h},a_{t:t+h},r_{t:t+h})$ from $\mathcal{E}$.
4:Sample a predicted action plan: $a^{\prime}_{t:t+h}\sim\pi(x_{t})$.